OpenStudy (anonymous):

I just need someone to set it up for me. A blueprint for a house states that 2.5 inches equals 10 feet. If the length of a wall is 12 feet, how long is the wall in the blueprint? a. 1 in. c. 3 in. b. 2 in. d. 4 in.

OpenStudy (ash2326):

@kittles<3 So we have the scale given, in the blue print 2.5 inches length is equal to 10 feet length in the house. We need to find the length of a wall in blue print which is 12 feet in length. Can you try?

OpenStudy (anonymous):

its cross multiply?

OpenStudy (ash2326):

yes, we can do that way as well, let x be the height of wall in blue print. And try to frame the equation now

OpenStudy (anonymous):

well can you set the problem up?

OpenStudy (ash2326):

It's easy to set up, 2.5 inch in blue print is 10 feet length in house. x feet in inch is 12 feet in length Both the ratios are same and therefore equal. You can set up the equation from here, can;t you?

OpenStudy (anonymous):

no no.I mean like 2.5/10=x/12 is that right?

OpenStudy (ash2326):

yes, it's right

OpenStudy (anonymous):

ohkay. (2.5)(x)=(10)(12) 2.5=120

OpenStudy (anonymous):

2.5x=120. 120/2.5=48

OpenStudy (ash2326):

you made a mistake \[\frac{2.5}{10}=\frac{x}{12}\]